If three people invest in a business and one puts in 300, one puts in 900, and one puts in 700, how do you divide up 95 in profit based on there investment?
Answer by stanbon(48545)

If three people invest in a business and one puts in 300, one puts in 900, and one puts in 700, how do you divide up 95 in profit based on there investment?
---------------------------------------------
The three people invested at a ratio of 3:9;7
This is the same as 3x:9x:7x
-------------------------------
EQUATION:
3x+9x+7x = 95
19x = 95
x = 95/19
----------------
3x = 3(95/19) = 285/19 = $15
9x = 9(95/19) = 3*15 = $45
7x = 7(95/19) = 7*5 = $35
-----------------------------
The 300 investor gets $15
The $900 investor gets $45
The $700 investeor gets $35
==============================
Cheers,
Stan H.